Ace Car Rental advertises that a rental car costs $25 per day plus a charge of 10 cents per mile. For the same car, Better Car Rental advertises a price of $40 per day plus 5 cents per mile. For what number of miles is the cost of renting a car the same at both companies?
Answer by checkley79(3341)

25+.10X=40+.05X
.10X-.05X=40-25
.05X=15
X=15/.05
X=300 MILE WILL RESULT I THE SAME COST.
PROOF:
25+.10*300=40+.05*300
25+30=40+15
55=55